OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. If 3a - 9 = 6, what is the value of 4a - 1?

OpenStudy (anonymous):

3a-9=6 3a=15 a=5 Plug in: 4a-1=? 4(5)-1=? 20-1=19 19

OpenStudy (agreene):

solve: 3a-9=6 then plug a in for : 4a-1 solving: 3a-9=6 add 9 to both sides 3a=15 divide both sides by 3 a=5 plug in 4(5)-1 20-1 19

OpenStudy (agreene):

i suppose, you should plug back in, since it says to: 3(5)-9=6 15-9=6 6=6 it checks